[figure: 9.7 cm]

XPS AND MAGNETIC STUDIES OF SmNi_{4}B COMPOUND

T. Toliński^{a}, G. Chełkowska^{b}, B. Andrzejewski^{a}, A. Kowalczyk^{a}, M. Timko^{c}, J. Kovac^{c}

aInstitute of Molecular Physics, PAS, Smoluchowskiego 17, 60-179 Poznań, Poland bInstitute of Physics, Silesian University, Uniwersytecka 4, 40-007 Katowice, Poland cInstitute of Experimental Physics, SAS, Watsonowa 47, 043 33 Košice, Slovakia

Our investigations show that the SmNi_{4}B compound is ferromagnetic with the ordering temperature T_{C} = 39 K obtained from the temperature dependence of magnetization. Our measurements of the magnetic moment versus the external magnetic field yield an evidence of a very large coercive field, H_{C} > 7.0 T. The magnetic moment μ(SmNi_{4}B) equal to 0.32 μ_{B}/f.u. in H = 9 T is reduced in comparison with the free ion value (μ(Sm) = 0.71 μ_{B}). The large H_{C} value is indicative of a significant anisotropy of this compound. X-ray photoemission spectroscopy shows that the valence band is well separated from the 4f peak. Between the Ni 2p_{1/2} and Ni 2p_{3/2} bands a satellite is visible, which may be related to the existence of a small magnetic moment on the Ni atoms. The measured hysteresis loop reveals no saturation of the magnetic moment in the applied magnetic field of 9 T.

[figure: 13.4 cm]

Subject category :

4. Rare Earthes and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : T. Toliński

Address for correspondence : Institute of Molecular Physics, PAS Smoluchowskiego 17, 60-179 Poznań Poland

Email address : tomtol@ifmpan.poznan.pl